Suppose a is rational and b is irrational. Show that a+b is irrational.
1
Expert's answer
2019-11-13T11:30:32-0500

Assume that "a+b=r" is a rational number. Since "r" and "a" are rational numbers, we obtain that "r-a" is a rational number. We obtain a contradiction, because "b=r-a" is not a rational number.

So our assumption that "a+b\in\mathbb Q" is false, that is "a+b".is a irrational number.